I'm not able to find the y-intercept in 3x+y=4. Some one please help!!


Thank you

Answer by ptaylor(1894)

I need to find the y-intercept of 3x+y=4.
This is the equation of a straight line. The standard form for this equation is :
y=mx+b
Where m is the slope and b is the y-intercept. This is called the slope- intercept form for a straight line.
Now we need to put your equation into standard form:
3x+y=4 subtract 3x from both sides
3x-3x+y=-3x+4
y=-3x+4 This is now in standard slope-intercept form and we can see that the y-intercept is +4 --------------ans

Again, I sure hope this helps-----ptaylor